OpenStudy (anonymous):

intergral of square root of[1+(cosx)^2]

OpenStudy (anonymous):

what the................

OpenStudy (luigi0210):

This? \(\Large \int \sqrt{1+(cosx)^2}~dx\)?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

you can do a change of variable and take X=cos(x) and instead of integrating over IR you will integrate over [-1,1]

OpenStudy (anonymous):

it's just an idea let's see if it will work

OpenStudy (anonymous):

let's replace dx first using X dX=-sin(x)dx let's replace again sin(x) by an expression using X so we know that (sin(x))^2+(cos(x))^2=1

OpenStudy (anonymous):

we obtain then \[\sin(x)=\sqrt{1-\cos^2(x)} \] we replace cos(x) by X then \[\sin(x)=\sqrt{1-X}\]

OpenStudy (anonymous):

sorry it's \[\sin(x)=\sqrt{1-X^2}\]

OpenStudy (anonymous):

wow thanks. I was stuck for the past 30min.]

OpenStudy (anonymous):

you're welcome except I don't know where it's goinbg to lead it's just an idea ;)

OpenStudy (loser66):

I am sorry for my stupid question. I don't get!! since the original problem is \(\Large \int \sqrt{1+(cosx)^2}~dx\), not \(\Large \int \sqrt{1-(cosx)^2}~dx\) but after manipulate the expression , you get the latter one, how to convert to the original one?

OpenStudy (anonymous):

I am computing the length of a curve where the original fn is y=sinx on the interval [0;pi]

OpenStudy (xapproachesinfinity):

@Loser66 that's just dX=-sinxdx=-root(1-X^2)dx so you have dx=dX/-root(1-X^2)

OpenStudy (xapproachesinfinity):

according to what he did applying change of variable

OpenStudy (xapproachesinfinity):